OpenStudy (anonymous):

The domain of the parabola x - 2 = -4(y - 3)^2 is

OpenStudy (anonymous):

is x > 2 right?

OpenStudy (anonymous):

how would i even graph this on a calculater considering its (y-3)^2? i always get confused!

OpenStudy (anonymous):

the coefficient on the y side is negative, so id go with x<2

OpenStudy (anonymous):

oh yes good point! didnt notice that, do you know how i could isolate y in an equation lik ethis?

OpenStudy (anonymous):

x-2 = -4(y-3)^2 (x-2)/-4 = (y-3)^2 \[\sqrt{(x-2)/-4} = y-3\] \[\sqrt{(x-2)/-4} +3 = y\]

OpenStudy (anonymous):

@wom u are a lifesaver!!! Thanks a bunch!

OpenStudy (anonymous):

wombat* lol

OpenStudy (anonymous):

Sure thing :) oh and the side with sqrt((x-2)/-4) +3 is going to have a + and - in front of it i think